OpenStudy (anonymous):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (25, 30) and (125, 50), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points? (Write your answer in simplest form, using / for a fraction bar if needed.)

OpenStudy (anonymous):

go for it Sam!

sam (.sam.):

 125-25=100 50-30=20 Slope=\[\sqrt{100^2+20^2}=101.98\]